Show your solution

A potential difference of 10.0 volts exists between two points, A and B, within an electric field. What is the magnitude of charge that requires 2.0 × 10–2 Joule of work to move it from A to B? (V = W / q


1
Expert's answer
2021-02-28T07:30:09-0500

Let's write the relationship between the electrical potential difference, work done and the amount of charge:


"\Delta V=V_B-V_A=\dfrac{W}{q},""q=\dfrac{W}{\Delta V}=\dfrac{2.0\cdot10^{-2}\ J}{10\ V}=2\cdot10^{-3}\ C."

Answer:

"q=2\cdot10^{-3}\ C."